# Correlation Properties in Channels with von Mises-Fisher Distribution of Scatterers

### v1

@@ Abstract
This letter presents simple analytical expressions for the spatial and temporal correlation functions in channels with von Mises-Fisher scattering. In contrast to previous results, the expressions presented here are exact and based only on elementary functions, clearly revealing the impact of the underlying parameters. The derived results are validated by a comparison against numerical integration result, where an exact match is observed. To demonstrate their utility, the presented results are used to analyze spatial correlation across different antenna array geometries and to investigate temporal correlation of a fluctuating radar signal from a moving target.
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication.
Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
The von Mises-Fisher (vMF) distribution is a widely adopted statistical model for angular distribution of scatterers in wireless communication channels, e.g., [1, 2], taking both the azimuth and elevation aspects of propagating waves into account. Its good fit to measurements and the ability to model non-isotropic scattering with different degrees of angular concentration of multipath components are its main advantages over the alternative models in the literature [3]. Due to its simplicity, the vMF distribution is preferred over more general models, such as the Fisher-Bingham (FB5) distribution [4].
While the vMF model is widely adopted, the associated statistics and the impact of the underlying parameters are not fully understood yet, primarily due to the lack of closed-form analytical results. The existing studies of spatial correlation characteristics are based on numerical integration [3], employ spherical wave expansion involving expressions with infinite sums of spherical Bessel functions [5], or obtain approximate analytical solutions under restrictive assumptions [6]. While these studies have greatly contributed to the initial understanding of non-isotropic scattering channels, the provided expressions are cumbersome and prohibit analytical inference, or their applicability is very restricted.
@@
The rest of this letter is structured as follows. Channel model is introduced in Section II, while Section III presents the derivation of the spatial and temporal correlation functions. In Section IV, the obtained results are validated and used to analyze spatial correlation across different antenna arrays and temporal correlation of a signal from a moving radar target. The paper is concluded in Section V.
@@ II Channel Model
The multipath components are assumed to arrive at the Rx with similar amplitudes, none of them being dominant. According to the central limit theorem, the envelope of (1) then exhibits Rayleigh distribution in this case. While this requires $N\rightarrow\infty$ in the theoretical reference model, it was demonstrated that $N\geq 10$ is practically sufficient [9].
@@ III-A Spatial Correlation Function
Spatial Correlation Function (SCF) of the complex baseband envelope, normalized to unit power, is obtained as [9]
@@
Employing [12, Eq. 3.338.4] to solve the inner integral in azimuth angle variable yields
@@
The SCF depends on both the displacement distance and the displacement direction relative to the mean DoA, with the dependence on the latter being controlled by the spread parameter $\kappa$. For $\kappa=0$, the dependence on the displacement direction disappears [^2] In the limit as $\kappa\rightarrow 0$, the normalization factor in front of the $\mathrm{sinc}$ function in (15) is equal to one, as can be shown by applying L’Hopital’s rule. Formally, a continuous extension of the function (15) is considered for $\kappa=0$. and the SCF reduces to the classical result for 3D isotropic scattering [13], i.e.
@@
The SCF is real in this case. However, the SCF in (15) is complex in general, with its real part corresponding to the auto-correlation of the quadrature components and the imaginary one to their cross-correlation, while its magnitude corresponds to the corelation of the complex envelope [9].
@@ III-B Auto-correlation Function
In addition to typical mobile communication scenarios with static scatterers considered in Section II, (19) also applies when the cluster of scatterers is the one being dynamic, with the Tx/Rx being static. This is the case in radar systems, where scattering from a moving target be represented by a vMF-distributed cluster of multipaths and (19) can be thereby used to analyze temporal correlation characteristics of the corresponding signal fluctuations at the radar Rx.
@@ III-C Multi-cluster Scattering
The results in (15) and (19) apply for a single cluster of vMF-distributed scatterers. However, their generalization to the multi-cluster scattering case typical for practical scenarios is straightforward. As initially derived for 2D von Mises scattering in [15], the SCF in channels with multiple vMF scattering clusters can be written as [^3] Here we explicitly indicate the dependence on vMF scattering parameters, as they generally differ between clusters.
@@
normalized power in $k$-th cluster, i.e.,
@@
SCF for $k$-th cluster only, i.e., given by (15);
@@
spread parameter for $k$-th cluster;
@@
mean DoA unit vector for $k$-th cluster.
@@ IV-A Impact of vMF Distribution Parameters on SCF
To verify the correctness of the derived result in (15), Fig. 1 additionally shows SCF curves obtained via numerical integration (dotted lines). The exact match is observed in all cases, thereby validating the presented results.
@@ IV-B Spatial Correlation over Antenna Arrays
As observed in Fig. 2a for $\kappa=10$, concentrated scattering results in a higher correlation over the planar array in the direction of the mean DoA projected onto the array surface. More concentrated scattering (higher $\kappa$) results in a wider and more elongated area of increased spatial correlation in direction of the mean DoA. On the other hand, for isotropic scattering ($\kappa=0$), the SCF is radially symmetric and exhibits variations according to the $\operatorname{sinc}$ function (16) in all directions.
@@ IV-C Temporal correlation of a fluctuating radar target
Fig. 3 shows the ACF of the Rx signal at a monostatic radar received for a target at an elevation angle of 20, represented by a vMF-distributed cluster of scatterers, with size specified by its angular width, $\Delta\theta$. The velocity vector of the target is assumed to be constant, horizontal and pointing in direction away from the radar. The operating frequency of the radar is assumed to be 10 GHz. The choice of the concentration parameter $\kappa$ for a given $\Delta\theta$ is detailed in Appendix A. To circumvent the numerical instability issues described in Appendix B, the approximation (23) is used to evaluate the SCF for the large values of $\kappa$ in this scenario.
@@
It is important to point out that the direction of target motion also has a significant impact on the Rx signal de-correlation which is not considered here, but can be anticipated based on in Fig. 1b. The Rx signal de-correlation with the change in the target’s orientation is neglected here, but has an important impact in practical radar systems. Similarly, the frequency correlation needs to be appropriately modelled in systems employing frequency agility techniques [14].
@@ V Conclusions
The vMF distribution is a widely adopted model for representing clusters of scatterers in wireless communication channels, taking both the main direction and concentration of scattering into account. This letter presents a simple closed-form expressions for the SCF and ACF in channels with vMF scattering. The simplicity of the presented expressions provides a better insight into the impact of the vMF distribution parameters on spatial correlation properties, and allows for analysis of arbitrary antenna array structures.

### v2

@@ Abstract
This letter presents simple analytical expressions for the spatial and temporal correlation functions in channels with von Mises-Fisher scattering. In contrast to previous results, the expressions presented here are exact and based only on elementary functions, clearly revealing the impact of the underlying scattering parameters. The derived results are validated by a comparison against numerical integration, where an exact match is observed. To demonstrate their utility, the presented results are used to analyze spatial correlation across different antenna array geometries and to investigate the temporal correlation of a fluctuating radar signal from a moving target.
@@
publicationid: pubid: ©2024 IEEE. Personal use of this material is permitted.
Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
The von Mises-Fisher (vMF) distribution is a widely adopted statistical model for angular distribution of scatterers in wireless communication channels, e.g., [1, 2], taking both the azimuth and elevation aspects of propagating waves into account. The good fit to measurements and the ability to model non-isotropic scattering with different degrees of angular concentration of multipath components are its main advantages over the alternative models in the literature [3]. Due to its simplicity, the vMF distribution is preferred over more general models, such as the Fisher-Bingham (FB5) distribution [4].
While the vMF model is widely adopted, the associated statistics and the impact of the underlying parameters are not fully understood yet, primarily due to the lack of closed-form analytical results. The existing studies of spatial correlation characteristics are either based on numerical integration [3], employ spherical wave expansion involving expressions with infinite sums of spherical Bessel functions [5], or obtain approximate analytical solutions under restrictive assumptions [6]. While these studies have greatly contributed to the initial understanding of non-isotropic scattering channels, the provided expressions are cumbersome and prohibit analytical inference, or their applicability is very restricted.
@@
The rest of this letter is structured as follows. Channel model is introduced in Section II, while Section III presents the derivation of the spatial and temporal correlation functions. In Section IV, the obtained results are validated and used to analyze spatial correlation across different antenna arrays and to investigate temporal correlation of a signal from a moving radar target. The paper is concluded in Section V.
@@ II Channel Model
The multipath components are assumed to arrive at the Rx with similar amplitudes, none of them being dominant. According to the central limit theorem, the envelope of (1) exhibits Rayleigh distribution in this case. While this requires $N\rightarrow\infty$ in the theoretical reference model, it was demonstrated that $N\geq 10$ is practically sufficient [9].
@@ III-A Spatial Correlation Function
Spatial Correlation Function (SCF) of the complex baseband envelope, normalized to unit power, is obtained as [^2] We should note that the SCF is frequency-dependent, as imposed by the presence of the wavelength $\lambda$ in (5). [9]
@@
Employing [12, Eq. 3.338.4] to solve the inner integral in the azimuth angle variable yields
@@
As we observe, the SCF depends on both the displacement distance and the displacement direction relative to the mean DoA, with the dependence on the latter being controlled by the spread parameter $\kappa$. For $\kappa=0$, the dependence on the displacement direction disappears [^3] In the limit as $\kappa\rightarrow 0$, the normalization factor in front of the $\mathrm{sinc}$ function in (15) is equal to one, as can be shown by applying L’Hopital’s rule. Formally, a continuous extension of the function (15) is considered for $\kappa=0$. and the SCF reduces to the classical result for 3D isotropic scattering [13], i.e.
@@
The SCF is real in this case. However, the SCF in (15) is complex in general, with its real part corresponding to the auto-correlation of the quadrature components and the imaginary one to their cross-correlation [9].
@@ III-B Auto-correlation Function
In addition to typical mobile communication scenarios with static scatterers considered in Section II, (19) also applies when the scattering cluster is the one being dynamic, with the Tx/Rx being static. This is the case in radar systems, where scattering from a moving target be represented by a vMF-distributed cluster of multipaths and (19) can be thereby used to analyze temporal correlation characteristics of the corresponding signal fluctuations at the radar Rx.
@@ III-C Multi-cluster Scattering
The results in (15) and (19) apply for a single cluster of vMF-distributed scatterers. However, their generalization to the multi-cluster scattering case typical for practical scenarios is straightforward. As initially derived for 2D von Mises scattering in [15], the SCF in channels with multiple vMF scattering clusters can be written as [^4] Here we explicitly indicate the dependence on vMF scattering parameters, as they generally differ between clusters. [3]
@@
normalized power in the $k$-th cluster, i.e.,
@@
SCF for the $k$-th cluster only, i.e., given by (15);
@@
spread parameter for the $k$-th cluster;
@@
mean DoA unit vector for the $k$-th cluster.
@@ IV-A Impact of vMF Distribution Parameters on SCF
To verify the correctness of the derived result in (15), Fig. 1 additionally shows SCF curves obtained via numerical integration (dotted lines). The exact match is observed in all cases, thereby validating the analytically derived results.
@@ IV-B Spatial Correlation over Antenna Arrays
As observed in Fig. 2a for $\kappa=10$, concentrated scattering results in a higher correlation over the planar array in the direction of the mean DoA projected onto the array surface. More concentrated scattering (higher $\kappa$) results in a wider and more elongated area of increased spatial correlation in the direction of the mean DoA. On the other hand, for isotropic scattering ($\kappa=0$), the SCF is radially symmetric and exhibits variations according to the $\operatorname{sinc}$ function (16) in all directions.
@@ IV-C Temporal correlation of a fluctuating radar target
Fig. 3 shows the ACF of the Rx signal at a monostatic radar received for a target at an elevation angle of 20, represented by a vMF-distributed cluster of scatterers, with size specified by its angular width, $\Delta\theta$. The velocity vector of the target is assumed to be constant, horizontal and pointing in the direction away from the radar. The operating frequency of the radar is assumed to be 10 GHz. The choice of the concentration parameter $\kappa$ for a given $\Delta\theta$ is detailed in Appendix A. To circumvent the numerical instability issues described in Appendix B, the approximation (23) is used to evaluate the SCF for the large values of $\kappa$ in this scenario.
@@
It is important to point out that the direction of target motion also has a significant impact on the Rx signal de-correlation which is not considered here, but can be anticipated based on the results in Fig. 1b. The Rx signal de-correlation with the change in the target’s orientation is neglected here, but has an important impact in practical radar systems. Similarly, the frequency correlation needs to be appropriately modelled in systems employing frequency agility techniques [14].
@@ V Conclusions
The vMF distribution is a widely adopted model for representing clusters of scatterers in wireless communication channels, taking both the main direction and the concentration of scattering into account. This letter presents a simple closed-form expressions for the SCF and ACF in channels with vMF scattering. The simplicity of the presented expressions provides a better insight into the impact of the vMF distribution parameters on spatial correlation properties, and allows for analysis of arbitrary antenna array structures.